# On nonlinear transitions, minimal seeds and exact solutions for the geodynamo

### v1

@@ Abstract
Nearly fifty years ago, Roberts (1978) postulated that Earth’s magnetic field, which is generated by turbulent motions of liquid metal in its outer core, likely results from a subcritical dynamo instability characterised by a dominant balance between Coriolis, pressure and Lorentz forces. Here we numerically explore the generation of subcritical geomagnetic fields using techniques from optimal control and dynamical systems theory to uncover the nonlinear dynamical landscape underlying dynamo action. Through nonlinear optimisation, via direct-adjoint looping, we identify the minimal seed — the smallest magnetic field that attracts to a nonlinear dynamo solution. Additionally, using the Newton-hookstep algorithm, we converge stable and unstable travelling wave solutions to the governing equations. By combining these two techniques, complex nonlinear pathways between attracting states are revealed, providing insight into a potential subcritical origin of the geodynamo. This paper showcases these methods on the widely studied benchmark of Christensen et al. (2001), laying the foundations for future studies in more extreme and realistic parameter regimes. We show that the minimal seed reaches a nonlinear dynamo solution by first attracting to an unstable travelling wave solution, which acts as an edge state separating a hydrodynamic solution from a magnetohydrodynamic one. Furthermore, by carefully examining the choice of cost functional, we establish a robust optimisation procedure that can systematically locate dynamo solutions on short time horizons with no prior knowledge of its structure.
@@ 1 Introduction
1.1 Balances in the geodynamo
@@
The problem of generation of Earth’s magnetic field is an important and challenging one for geophysics; the magnetic field plays a key role in deflecting the solar wind and therefore is important in helping to protect us from the harmful effects of the charged particles and the depletion of ozone leading to increased exposure to harmful ultraviolet radiation (see e.g. Arsenović et al., 2024). The geomagnetic field is predominantly dipolar, and so large scale, with a dipole axis that is offset from the axis of rotation. It has variability on a wide range of timescales from years to centuries and includes such dynamical features as excursions and reversals (see e.g. Constable and Korte, 2006; Roberts, 2008).
@@
Earth’s magnetic field is believed to be generated by dynamo action in its fluid outer core where convective motions (either compositionally or thermally driven) in the electrically conducting medium can drive currents and hence magnetic fields via induction (see e.g. Moffatt and Dormy, 2019). The geodynamo problem involves the simultaneous solution of the partial differential equations governing the evolution of the fluid (the momentum equation), the magnetic field (the induction equation) and the evolution equation for the co-density. The theoretical challenge arises owing to the vast range of temporal scales (and to a lesser extent spatial scales) that need to be resolved in the momentum equation (with the induction and codensity equations causing fewer problems); the large range of temporal scales leads to the requirement that the non-dimensional equations be solved at extreme parameter values and hence the impossibility of direct solution via computational methods.
@@
Nonetheless progress has been made on a number of fronts. Heroic computations (following on from the pioneering work of Glatzmaier and Roberts, 1995) are starting to achieve small enough parameters (though still extremely far from their true values) that interesting dynamics are achieved (see e.g. Aubert et al., 2017; Schaeffer et al., 2017; Schwaiger et al., 2019). This computational endeavour has been predicated on the understanding that it is important to approach the correct regime by keeping the balances in the momentum equation and codensity equations realistic as one increases the separation of timescales. This is achieved by considering a distinguished limit, as suggested by Dormy (2016), so that all the non-dimensional parameters scale together to achieve the correct balance.
@@
What is the correct balance for the geodynamo? It is widely thought that the magnetic field in the Earth’s core is strong enough to play a significant role in the dynamics of the flow; at least at some lengthscale. For a detailed discussion of how, and at what scale, the magnetic Lorentz force may enter in the dynamics, see Aurnou and King (2017) or Tobias (2021). Theoretical support for the importance of the magnetic field arises from theoretical and computational studies of magnetoconvection (see e.g. Proctor, 1994; Fearn, 1998). Roberts (1978) recognised that there may be two branches of the Earth’s dynamo; a weak field branch and a strong field branch. The weak field branch arises in a traditional bifurcation scenario; as the thermal driving (as measured by the non-dimensional Rayleigh number) is increased hydrodynamic thermal convection sets in. Because the Earth is a rapid rotator, this convection sets in at very small length-scale and at very high Rayleigh number. Eventually the convection becomes strong enough so that the magnetic Reynolds number of the flow is large enough for dynamo action to onset. This leads to the formation of weak multipolar magnetic fields (Petitdemange, 2018); here these fields are weak in the sense that their magnetic energy is equivalent to, or less than, the kinetic energy of the flow.
@@
A more efficient situation occurs when the magnetic field is strong. The strong magnetic field breaks the restriction imposed on the flow by the fast rotation and leads to the onset of convection on larger lengthscales at more moderate values of the Rayleigh number measuring thermal driving. Thus one might expect to find a more efficient subcritical strong field branch for values of the Rayleigh number well below onset of weak field dynamo action (and potentially even below the onset of hydrodynamic convection). On the strong field branch the magnetic energy should be much larger than the kinetic energy of the flow and the Lorentz force may even be strong enough to enter into the leading order balance, competing with the Coriolis force and the pressure gradient to give a magnetostrophic balance (at least at certain scales (Dormy, 2016)). The magnetic field in this situation acts so as to make the dynamo behave in a more efficient manner — an example of an essentially nonlinear dynamo (Tobias et al., 2011). There is therefore a strong theoretical drive to understand subcritical dynamo action in the context of rapidly rotating spherical shells.
1.2 Essentially nonlinear dynamos
@@
It is, however, non-trivial to compute nonlinear dynamos in the subcritical regime. Although timestepping methods are able sometimes to locate these, the results are often haphazard and sometimes owe their success to a slice of luck. A more systematic approach relies on the computation of fixed points. The latter builds on quasi-Newton iterative solvers to compute equilibrium dynamo solutions by continuation while progressing through the parameter space. Such solvers however require fairly good first guesses for the procedure to converge toward a solution. This usually implies finding a starting point in the linearly unstable regime (whenever possible) and then following the saturated solution down as the control parameter is gradually decreased toward the linearly stable regime. Alternatively, one may turn the subcritical problem into a supercritical one e.g. through the addition of a suitable physical forcing, which is then gradually removed (Waleffe, 2003) — in a dynamo context, this method was successfully used by Rincon et al. (2007) to identify subcritical equilibrium solutions in rotating, magnetohydrodynamic (MHD) plane Couette flow. Moreover, while this continuation approach has the considerable advantage of computing unstable as well as stable equilibria, it can only compute exact solutions that are travelling waves or at least recurrent states, which limits its applicability.
@@
To overcome these limitations, a robust way to numerically identify an elusive subcritical state without knowing much about it is through a combination of short-time optimal control and long-time direct numerical simulations, as demonstrated by Mannix et al. (2022) with nonlinear dynamos. This alternative approach builds on the mathematical framework of transient growth and nonmodal stability analysis (Schmid, 2007), which has been widely used in the last decade to compute ‘minimal’ perturbations that nonlinearly trigger transition to hydrodynamic turbulence in shear flows (Pringle and Kerswell, 2010; Kerswell, 2018). Note that in the context of sheared turbulence, the objective is not to find the subcritical state, which (in strong contrast to subcritical dynamo problems) is fairly easy to trigger numerically. Instead, the objective is rather to identify the smallest-energy perturbation that kickstarts turbulence, which no alternative methods can currently compute. In particular, fixed point methods cannot identify minimal seeds since these are not steady or recurring solutions of the underlying dynamical system. Mannix et al. (2022) found that the spatial structure of minimal dynamo seeds bears no resemblance to that of the subcritical, nonlinear state — a finding that is consistent with studies of minimally-triggered turbulence in shear flows (Pringle and Kerswell, 2010), and the sequence of events they trigger is revealing of the mechanisms of subcritical dynamo action in the considered systems.
@@
In this paper, we investigate the geodynamo benchmark of Christensen et al. (2001) as an ideal starting point for developing the methodology needed to identify minimal seeds that lead to subcritical dynamo action in a geodynamo setup. This benchmark study contains three cases, of which we consider the first two; a purely hydrodynamic benchmark and a subcritical dynamo benchmark. In the hydrodynamic benchmark, convective instability gives rise to Taylor columns with $m=4$ symmetry that ultimately persists as a travelling wave. Although at the benchmark parameters the conducting (steady base-state) is unstable to convection, the resulting convecting state is subcritically stable to dynamo action. In this manner, the form of the magnetic field initial condition is crucial, and needs to have a large enough amplitude as well as the correct structure in order to reach a nonlinear dynamo.
@@
Although originally intended as a benchmark for Boussinesq MHD spherical shell codes, the ‘modest’ benchmark parameters provide rich dynamical behaviour that has subsequently been studied. The hydrodynamic bifurcation behaviour has been explored by Feudel et al. (2015), which shows that there are multiple stable nonlinear hydrodynamic travelling waves present at the benchmark parameters. Subsequently, Feudel et al. (2017) studied the bifurcation nature of nonlinear dynamo solutions possible in the system, showing that at the benchmark parameters there are stable and unstable MHD states that arise via a saddle node bifurcation, and that the MHD solution branch is smoothly connected to a purely hydrodynamic branch where a secondary Hopf bifurcation occurs. In these studies, the nonlinear travelling waves were converged using a matrix-free Newton solver using implicit integration of the linear terms as a preconditioner (Mamun and Tuckerman, 1995). More recently, Skene and Tobias (2024) used a weakly nonlinear analysis to study the saturation of the convective instability in the purely hydrodynamic case near the onset of convection.
@@
While these studies illuminate the dynamical states possible for the system, they do not provide information on the complex nonlinear pathways between these states or their basins of attraction. A more systematic programme for mapping out the landscape of subcritical dynamos is needed, with tools developed for uncovering dynamics in a spherical shell. It is the first steps of just such a programme that we shall describe in this paper, which combines nonlinear optimisation, together with identification of the unstable and stable travelling waves present in the system. Taking techniques from optimal control and nonlinear optimisation, we can identify the magnetic field with the smallest energy that attracts to a dynamo solution; providing a limit on the basis of attraction for reaching a nonlinear dynamo, as well as systematically finding a dynamo solution without prior knowledge of its structure. By carefully considering the cost functional to extremise, we will furnish a robust technique that can identify dynamos using short time horizon optimisations, enabling this procedure to be scaled up in future studies. Using a Newton-hookstep procedure (Dennis and Schnabel, 1996), we will find stable and unstable travelling waves in the system, and examine them in light of the dynamical pathways between important states. We demonstrate that the Newton-hookstep procedure is able to converge unstable states without preconditioning, and therefore provide the foundation for future studies where these exact solutions will lie within a turbulent and chaotic flow field. The structure of the paper is as follows; the mathematical and numerical setup is outlined in §2, the results are detailed in §3, and conclusions are offered in section §4.
@@ 2.1 Governing equations
$\begin{gathered}\frac{\partial\hat{\bm{U}}}{\partial\hat{t}}+\hat{\bm{U}}\cdot\nabla\hat{\bm{U}}=-\frac{1}{\rho_{0}}\nabla\hat{P}-2\Omega\bm{e}_{z}\times\hat{\bm{U}}+\frac{g_{0}\alpha\hat{r}}{r_{0}}\hat{T}\bm{e}_{r}+\nu\nabla^{2}\hat{\bm{U}}+\frac{1}{\mu}(\nabla\times\bm{B})\times\bm{B},\\ \nabla\cdot\hat{\bm{U}}=0,\\ \frac{\partial\hat{T}}{\partial\hat{t}}+\hat{\bm{U}}\cdot\nabla\hat{T}=\kappa\nabla^{2}\hat{T},\end{gathered}$ (1)
along with the induction equation
$\begin{gathered}\frac{\partial\hat{\bm{B}}}{\partial\hat{t}}=\nabla\times(\hat{\bm{U}}\times\hat{\bm{B}})+\eta\nabla^{2}\hat{\bm{B}},\\ \nabla\cdot\hat{\bm{B}}=0.\end{gathered}$ (2)
These dimensional equations describe the evolution of the velocity field $\hat{\bm{U}}$, temperature field $\hat{T}$, magnetic field $\hat{\bm{B{}}}$ and modified pressure $\hat{P}$ (which also accounts for the centrifugal acceleration). Although we use the temperature, we note that one could also formulate equations (1) in terms of the codensity. In writing the equations in dimensional form, denoted with ‘hats’, we have used the kinematic viscosity $\nu$, the magnetic permeability $\mu$, the magnetic diffusivity $\eta$, the thermal expansion coefficient $\alpha$, the gravity at the outer radius $g_{o}$, and the thermal diffusivity $\kappa$. The outer region $\mathcal{V}_{o}$ is taken to be an insulator, and so the electrical current vanishes there: $\hat{\bm{J}}=\nabla\times\hat{\bm{B}}=0$, so that the magnetic field satisfies a Laplace equation in $\mathcal{V}_{o}$. Whilst different approaches have been used to model the inner region $\mathcal{V}_{i}$, such as taking it to be finitely conducting (Glatzmaier and Roberts, 1995), we will take it to be insulating.
@@
where we have introduced the notation $\bm{r}=r\bm{e}_{r}$. These equations are governed by the Ekman number Ek, the modified Rayleigh number ${\tilde{\textit{Ra}}}$, the magnetic Prandtl number Pm and the Prandtl number Pr, respectively defined as
@@
2.2 The optimisation procedure
For simplicity we will now write equations (3) in the general form
$\mathbf{M}\frac{\partial\bm{q}}{\partial t}=\bm{\mathcal{N}}(\bm{q}),$ (10)
where the state $\bm{q}=(\bm{U},T,P,\bm{B})^{T}$, $\mathbf{M}$ is chosen to ensure that the time derivatives do not apply to the divergence free conditions, and $\bm{\mathcal{N}}$ represents the non-linear right hand side of the equations. To derive the adjoint equations we first setup our variational problem as maximising a cost functional $\mathcal{J}$ subject to $\bm{q}$ satisfying the governing equations. This can be achieved by defining the Lagrangian
$\mathcal{L}=\mathcal{J}-\left[\bm{q}^{\dagger},\mathbf{M}\frac{\partial\bm{q}}{\partial t}-\bm{\mathcal{N}}(\bm{q})\right],$ (11)

### v2

@@ Abstract
Nearly fifty years ago, Roberts, (1978) postulated that Earth’s magnetic field, which is generated by turbulent motions of liquid metal in its outer core, likely results from a subcritical (finite-amplitude) dynamo instability characterised by a dominant balance between Coriolis, pressure and Lorentz forces. Here we numerically explore subcritical convective dynamo action in a spherical shell, using techniques from optimal control and dynamical systems theory to uncover the nonlinear dynamics of magnetic field generation. Through nonlinear optimisation, via direct-adjoint looping, we identify the minimal seed — the smallest magnetic field that attracts to a nonlinear dynamo solution. Additionally, using the Newton-hookstep algorithm, we converge stable and unstable travelling wave solutions to the governing equations. By combining these two techniques, complex nonlinear pathways between attracting states are revealed, providing insight into a potential subcritical origin of the geodynamo. This paper showcases these methods on the widely studied benchmark of Christensen et al., (2001), laying the foundations for future studies in more extreme and realistic parameter regimes. We show that the minimal seed reaches a nonlinear dynamo solution by first approaching an unstable travelling wave solution, which acts as an edge state separating a hydrodynamic solution from a magnetohydrodynamic one. Furthermore, by carefully examining the choice of cost functional, we establish a robust optimisation procedure that can systematically locate dynamo solutions on short time horizons with no prior knowledge of its structure.
@@ 1 Introduction
1.1 Balances in the geodynamo: weak and strong field branches
@@
The problem of generation of Earth’s magnetic field is an important and challenging one for geophysics; the magnetic field plays a key role in deflecting the solar wind and therefore is important in helping to protect us from the harmful effects of the charged particles and the depletion of ozone leading to increased exposure to harmful ultraviolet radiation (see e.g. Arsenović et al.,, 2024). The geomagnetic field is predominantly dipolar, and so large scale, with a dipole axis that is offset from the axis of rotation. It has variability on a wide range of timescales from years to centuries and includes such dynamical features as excursions and reversals (see e.g. Constable and Korte,, 2006; Roberts,, 2008).
@@
Earth’s magnetic field is believed to be generated by dynamo action in its fluid outer core where convective motions (either compositionally or thermally driven) in the electrically conducting medium can drive currents and hence magnetic fields via induction (see e.g. Moffatt and Dormy,, 2019). The geodynamo problem involves the simultaneous solution of the partial differential equations governing the evolution of the fluid (the momentum equation), the magnetic field (the induction equation) and the evolution equation for the co-density. The theoretical challenge arises owing to the vast range of temporal scales (and to a lesser extent spatial scales) that need to be resolved in the momentum equation (with the induction and codensity - a variable which combines the buoyancy effects of temperature and chemical species concentration (Braginsky and Roberts,, 1995) - equations causing fewer problems); the large range of temporal scales leads to the requirement that the non-dimensional equations be solved at extreme parameter values and hence the impossibility of direct solution via computational methods.
@@
Nonetheless progress has been made on a number of fronts. Heroic computations (following on from the pioneering work of Glatzmaier and Roberts,, 1995) are starting to achieve small enough parameters (though still extremely far from their true values) that interesting dynamics are achieved (see e.g. Aubert et al.,, 2017; Schaeffer et al.,, 2017; Schwaiger et al.,, 2019). This computational endeavour has been predicated on the understanding that it is important to approach the correct regime by keeping the balances in the momentum equation and codensity equations realistic as one increases the separation of timescales. This is achieved by considering a distinguished limit, as suggested by Dormy, (2016), so that all the non-dimensional parameters scale together to achieve the correct balance.
@@
What is the correct balance for the geodynamo? It is widely thought that the magnetic field in the Earth’s core is strong enough to play a significant role in the dynamics of the flow; at least at some lengthscale. For a detailed discussion of how, and at what scale, the magnetic Lorentz force may enter in the dynamics, see Aurnou and King, (2017) or Tobias, (2021). Theoretical support for the importance of the magnetic field arises from theoretical and computational studies of magnetoconvection (see e.g. Proctor,, 1994; Fearn,, 1998). Roberts, (1978) recognised that there may be two branches of the Earth’s dynamo; a weak field branch and a strong field branch. The weak field branch arises in a traditional bifurcation scenario; as the thermal driving (as measured by the non-dimensional Rayleigh number) is increased hydrodynamic thermal convection sets in. Because the Earth is a rapid rotator, this convection sets in at very small length-scale and at very high Rayleigh number. Eventually the convection becomes strong enough so that the magnetic Reynolds number of the flow is large enough for dynamo action to onset. This leads to the formation of weak multipolar magnetic fields (Petitdemange,, 2018); here these fields are weak in the sense that their magnetic energy is equivalent to, or less than, the kinetic energy of the flow.
@@
A more efficient situation occurs when the magnetic field is strong. The strong magnetic field breaks the restriction imposed on the flow by the fast rotation and leads to the onset of convection on larger lengthscales at more moderate values of the Rayleigh number measuring thermal driving. Thus one might expect to find a more efficient subcritical strong field branch for values of the Rayleigh number well below onset of weak field dynamo action (and potentially even below the onset of hydrodynamic convection). On the strong field branch the magnetic energy should be much larger than the kinetic energy of the flow and the Lorentz force may even be strong enough to enter into the leading order balance, competing with the Coriolis force and the pressure gradient to give a magnetostrophic balance (at least at certain scales (Dormy,, 2016)). The magnetic field in this situation acts so as to make the dynamo behave in a more efficient manner — an example of an essentially nonlinear dynamo (Tobias et al.,, 2011). There is therefore a strong theoretical drive to understand subcritical dynamo action in the context of rapidly rotating spherical shells.
1.2 Computing essentially nonlinear dynamos
@@
It is, however, non-trivial to compute nonlinear dynamos in the subcritical regime. Although timestepping methods are able sometimes to locate these, the results are often haphazard and sometimes owe their success to a slice of luck. A more systematic approach relies on the computation of fixed points. The latter builds on quasi-Newton iterative solvers to compute equilibrium dynamo solutions by continuation while progressing through the parameter space. Such solvers however require fairly good first guesses for the procedure to converge toward a solution. This usually implies finding a starting point in the linearly unstable regime (whenever possible) and then following the saturated solution down as the control parameter is gradually decreased toward the linearly stable regime. Alternatively, one may turn the subcritical problem into a supercritical one e.g. through the addition of a suitable physical forcing, which is then gradually removed (Waleffe,, 2003) — in a dynamo context, this method was successfully used by Rincon et al., (2007) to identify subcritical equilibrium solutions in rotating, magnetohydrodynamic (MHD) plane Couette flow. Moreover, while this continuation approach has the considerable advantage of computing unstable as well as stable equilibria, it can only compute exact solutions that are travelling waves or at least recurring states, which limits its applicability.
In the context of wall-bounded turbulence, the objective of such computations of equilibria is not to find the attracting, statistically steady turbulent state. This state is fairly easy to reproduce numerically (or indeed experimentally). The theoretical challenge rather lies with the identification of unstable coherent states and their dynamical role in the transition to turbulence (Toh and Itano,, 2003; Schneider et al.,, 2007; Wang et al.,, 2007; Khapko et al.,, 2016), with the control of the transition as a long-term objective. In the context of subcritical dynamo problems, however, identifying the attracting dynamo state in itself can be very uncertain. It is imperative that the computed dynamo states be allowed to be highly fluctuating. Furthermore, it is desirable that as little knowledge of the dynamo mechanism as possible be required for the computation. To overcome these limitations, a robust way to numerically identify an elusive subcritical state without prior guesses of the equilibrium is through a combination of short-time optimal control and long-time direct numerical simulations, as demonstrated by Mannix et al., (2022) with nonlinear dynamos.
@@
This alternative approach builds on the mathematical framework of transient growth and nonmodal stability analysis (see for example Schmid, (2007); Kerswell, (2018) for reviews of both linear and nonlinear nonmodal stability analysis, respectively). It has been widely used in the last decade to compute ‘minimal’ (i.e. least-energy) perturbations that nonlinearly trigger transition to hydrodynamic turbulence in shear flows (Pringle and Kerswell,, 2010; Duguet et al.,, 2010; Cherubini et al.,, 2010; Vavaliaris et al.,, 2020), which no alternative methods can currently compute. In particular, fixed point methods cannot identify minimal seeds since these are not steady or recurring solutions of the underlying dynamical system. The review of Kerswell et al., (2014) illustrates this approach for identifying minimal seeds in dynamical systems in general, and identifies three key considerations. The first outlines the importance of the optimisation time horizon, which must be large enough to overcome initial transients but not so large as to render convergence difficult due to the increasing sensitivity of the optimum to the initial condition. Secondly, including nonlinearities is imperative since the most-amplified perturbations in the linear problem are (generally) not the minimal seeds for subcritical transition. Lastly, the functional to be optimised to compute least-energy perturbations need not be the energy. Further to these points, Mannix et al., (2022) found that the spatial structure of minimal dynamo seeds bears no resemblance to that of the subcritical, nonlinear state – a finding that is consistent with studies of minimally-triggered turbulence in shear flows – and that the sequence of events they trigger is revealing of the mechanisms of subcritical dynamo action in the considered systems.
1.3 Triggering nonlinear transition to subcritical, convective dynamos
@@
In this paper, we investigate the geodynamo benchmark of Christensen et al., (2001) as an ideal starting point for developing the methodology needed to identify minimal seeds that lead to subcritical dynamo action in a geodynamo setup. This benchmark study contains three cases, of which we consider the first two; a purely hydrodynamic benchmark and a subcritical dynamo benchmark. In the hydrodynamic benchmark, convective instability gives rise to Taylor columns with $m=4$ symmetry that ultimately persists as a travelling wave. Although at the benchmark parameters the conducting (steady base-state) is unstable to convection, the resulting convecting state is linearly stable to dynamo action. In this manner, the form of the magnetic field initial condition is crucial, and needs to have a large enough amplitude as well as the correct structure in order to reach a nonlinear dynamo.
@@
Although originally intended as a benchmark for Boussinesq MHD spherical shell codes, the ‘modest’ benchmark parameters provide rich dynamical behaviour that has subsequently been studied. The hydrodynamic bifurcation behaviour has been explored by Feudel et al., (2015), which shows that there are multiple stable nonlinear hydrodynamic travelling waves present at the benchmark parameters. Subsequently, Feudel et al., (2017) studied the bifurcation nature of nonlinear dynamo solutions possible in the system, showing that at the benchmark parameters there are stable and unstable MHD states that arise via a saddle node bifurcation, and that the MHD solution branch is smoothly connected to a purely hydrodynamic branch where a secondary Hopf bifurcation occurs. In these studies, the nonlinear travelling waves were converged using a matrix-free Newton solver using implicit integration of the linear terms as a preconditioner (Mamun and Tuckerman,, 1995). More recently, Skene and Tobias, (2024) used a weakly nonlinear analysis to study the saturation of the convective instability in the purely hydrodynamic case near the onset of convection.
@@
While these studies illuminate the dynamical states possible for the system, they do not provide information on the complex nonlinear pathways between these states or their basins of attraction. A more systematic programme for mapping out the landscape of subcritical dynamos is needed, with tools developed for uncovering dynamics in a spherical shell. It is the first steps of just such a programme that we shall describe in this paper, which combines nonlinear optimisation with the identification of the unstable and stable travelling waves present in the system. Optimal control, on the one hand, can identify the magnetic field with the smallest energy that attracts to a dynamo solution. Variational optimisation in dynamo theory dates back to Backus, (1958), who derived a bound on the magnetic Reynolds number required for dynamo action in a sphere. More recently, optimisation of steady velocity fields has been performed in various geometries to maximise linear dynamo growth rates (Willis,, 2012; Chen et al.,, 2015, 2018; Herreman,, 2018). In the present paper we use optimal control on the full set of MHD equations to provide a limit on the basins of attraction for reaching a nonlinear dynamo. By considering the cost functional to extremise, we furnish a robust procedure that can identify dynamos using short time horizon optimisations, enabling this procedure to be scaled up in future studies. On the other hand, we also use a Newton-hookstep method (Dennis and Schnabel,, 1996) to find stable and unstable travelling waves in the system, and examine them in light of the dynamical pathways between important states. We demonstrate that the Newton-hookstep procedure is able to converge unstable magnetic states without preconditioning, and therefore provide the foundation for future studies where these exact solutions will lie within a turbulent and chaotic flow field. The structure of the paper is as follows; the mathematical and numerical setup is outlined in §2, the results are detailed in §3, and conclusions are offered in section §4.
@@ 2.1 Governing equations
$\begin{gathered}\frac{\partial\hat{\bm{U}}}{\partial\hat{t}}+\hat{\bm{U}}\cdot\hat{\nabla}\hat{\bm{U}}=-\frac{1}{\rho_{0}}\hat{\nabla}\hat{P}-2\Omega\bm{e}_{z}\times\hat{\bm{U}}+\frac{g_{0}\alpha\hat{r}}{r_{0}}\hat{T}\bm{e}_{r}+\nu\hat{\nabla}^{2}\hat{\bm{U}}+\frac{1}{\mu}(\hat{\nabla}\times\hat{\bm{B}})\times\hat{\bm{B}},\\ \nabla\cdot\hat{\bm{U}}=0,\\ \frac{\partial\hat{\bm{B}}}{\partial\hat{t}}=\hat{\nabla}\times(\hat{\bm{U}}\times\hat{\bm{B}})+\eta\hat{\nabla}^{2}\hat{\bm{B}},\\ \hat{\nabla}\cdot\hat{\bm{B}}=0,\\ \frac{\partial\hat{T}}{\partial\hat{t}}+\hat{\bm{U}}\cdot\hat{\nabla}\hat{T}=\kappa\hat{\nabla}^{2}\hat{T}.\end{gathered}$ (1)
These dimensional equations describe the evolution of the velocity field $\hat{\bm{U}}$, temperature field $\hat{T}$, magnetic field $\hat{\bm{B{}}}$ and modified pressure $\hat{P}$ (which also accounts for the centrifugal acceleration). Although we use the temperature, we note that one could also formulate equations (1) in terms of the codensity. In writing the equations in dimensional form, denoted with ‘hats’, we have used the kinematic viscosity $\nu$, the magnetic permeability $\mu$, the magnetic diffusivity $\eta$, the thermal expansion coefficient $\alpha$, the gravity at the outer radius $g_{o}$, and the thermal diffusivity $\kappa$. The outer region $\mathcal{V}_{o}$ is taken to be an insulator, and so the electrical current vanishes there: $\hat{\bm{J}}=\hat{\nabla}\times\hat{\bm{B}}=0$, so that the magnetic field satisfies a Laplace equation in $\mathcal{V}_{o}$. Whilst different approaches have been used to model the inner region $\mathcal{V}_{i}$, such as taking it to be finitely conducting (Glatzmaier and Roberts,, 1995), we will take it to be insulating.
@@
where we have introduced the notation $\bm{r}=r\bm{e}_{r}$. These equations are governed by the Ekman number Ek, the modified Rayleigh number ${\tilde{\textit{Ra}}}$ (which accounts for the stabilising effect of rotation on convection, and is related to the classical Rayleigh number $\textit{Ra}=\alpha g_{o}\Delta Td^{3}/(\nu\kappa)$ via $\tilde{\textit{Ra}}=\textit{Ra}\textit{Ek}/\textit{Pr}$), the magnetic Prandtl number Pm and the Prandtl number Pr, respectively defined as
@@
2.2 The adjoint-based optimal control procedure
We briefly outline the optimisation procedure for determining the initial condition $\bm{q_{0}}$ that maximises a chosen cost functional under the strong constraint that the system, described by the state variable $\bm{q}(\bm{x},t)$, evolves according to the governing equations, written here in general form:
$\mathbf{M}\frac{\partial\bm{q}}{\partial t}=\bm{\mathcal{N}}(\bm{q}),\qquad\bm{q}(\bm{x},0)=\bm{q_{0}}.$ (9)
This constrained optimisation problem can be solved by determining the critical points of an augmented Lagrangian
$\mathcal{L}(\bm{q},\bm{q_{0}},\bm{q}^{\dagger},\bm{q}^{\dagger}_{0})=\mathcal{J}(\bm{q},\bm{q_{0}})-\left[\bm{q}^{\dagger},\mathbf{M}\frac{\partial\bm{q}}{\partial t}-\bm{\mathcal{N}}(\bm{q})\right]-\langle\bm{q}^{\dagger}_{0},\bm{q}(\bm{x},0)-\bm{q_{0}}\rangle,$ (10)